1. If R1 is 2.3k ohms, R2 is 560 ohms & R3 is 24h ohms , what is Rt?

2. If V=4v , I=18uA, what is R1?
1
Expert's answer
2020-05-19T09:13:22-0400

If the resistors are connected in series, the total resistance equals the sum of all resistors, and it is


"R_{ts}=R_1+R_2+R_3=\\=2.3+0.56+0.024=2.884\text{ k}\Omega."

If the resistors are in parallel, the total resistance can be found as


"\frac{1}{R_{tp}}=\frac{1}{R_1}+\frac{1}{R_2}+\frac{1}{R_3},\\
R_{tp}=0.023\text{ k}\Omega."


If the voltage across the first resistor is 4 V and the current in the first resistor is 18 microamps, the resistance will be


"R_1=\frac{V}{I}=\frac{4}{18\cdot10^{-6}}=222.22\text{ k}\Omega."